OpenStudy (anonymous):

find the limit please \[\lim_{x \rightarrow infinity}\sqrt{x^2+3x+1}-x\]

myininaya (myininaya):

you should put a 1 underneath that and rationalize the numerator by multiplying top and bottom by the conjugate of the top

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{\sqrt{x^2+3x+1}-x}{1} \cdot \frac{\sqrt{x^2+3x+1}+x}{\sqrt{x^2+3x+1}+x}\]

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{(x^2+3x+1)-x^2}{\sqrt{x^2+3x+1}+x}\]

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{3x+1}{\sqrt{x^2+3x+1}+x} \cdot \frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}} \]

myininaya (myininaya):

|x|=x if x>0 which it is since x gets positive large

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{3+\frac{1}{x}}{\sqrt{1+\frac{3}{x}+\frac{1}{x^2}}+\frac{x}{x}}\]

myininaya (myininaya):

\[\frac{3+0}{\sqrt{1+0+0}+1}\]

OpenStudy (anonymous):

thanks once again

myininaya (myininaya):

lol np

OpenStudy (anonymous):

this is the last one, can you help me with it please

OpenStudy (anonymous):

Compute A= x^2+x/x+1 as x approaches +infinity and B= x^2+x/x^2+1 as x approaches -infinity

OpenStudy (anonymous):

\[\lim_{x \rightarrow +\infty} (x^2+x) / (x+1) = +\infty\]\[\lim_{x \rightarrow -\infty}(x^2 +x)/(x^2+1) =1\]